### v2

# Beyond the Electric Dipole Approximation: Electric and Magnetic Multipole Contributions Reveal Biaxial Water Structure from SFG Spectra at the Air-Water Interface

Abstract
The interpretation of sum-frequency generation (SFG) spectra has been severely limited by the absence of quantitative theoretical predictions of higher-order multipole contributions. The previously unknown magnetic dipole and electric quadrupole contributions are determined by bulk properties but appear in all experimental SFG spectra, obscuring the connection between measured spectra and interfacial structure. We present the simulation-based framework that predicts the full set of multipole spectral contributions. This framework also yields depth-resolved spectra, enabling the precise localization of spectroscopic features. Applied to the air-water interface, our approach achieves quantitative agreement with experimental spectra for different polarization combinations in both the bending and stretching regions. Higher-order multipole contributions are crucial for correctly interpreting SFG spectra: in the bending band, the electric dipole and the magnetic dipole contributions have similar intensities, and the electric quadrupole contribution is significantly larger. In the OH-stretch region, the electric quadrupole contribution is found to be responsible for the mysterious shoulder at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$. Crucially, subtracting the quadrupole and magnetic contributions isolates the second-order electric dipole susceptibility, which is a quantitative probe for interfacial orientational anisotropy. This electric-dipole susceptibility reveals a pronounced biaxial ordering of water at the air-water interface. By resolving a fundamental limitation of the interpretation of SFG spectroscopy, our framework allows for the detailed extraction of interfacial water ordering from SFG spectra.
I Introduction
Interfaces are crucial in biological [1, 2, 3, 4, 5] and physicochemical [6, 7, 8, 9, 10] applications. In particular, the air-water interface plays an essential role in catalysis [11, 12, 13, 14, 15], atmospheric [16, 17, 18], and prebiotic [19, 20] chemistry. Undoubtedly, understanding the microscopic structure of interfaces is essential. The inhomogeneous interfacial region is only a few angstroms thick [21], or up to a few nanometers if the interface is charged [22, 23], and thus much thinner than the bulk region. This presents significant challenges in spectroscopy, as signals from the interface must be distinguished from bulk signals. An elegant solution is to measure the second-order susceptibility, which is non-zero only when the spatial inversion symmetry is broken, using sum frequency generation (SFG) spectroscopy. Typically, the bulk medium is isotropic and, therefore, the interface’s structure accounts for a significant part of the SFG spectrum. The interpretation of experimental spectra usually assumes that second-order radiation originates from an interfacial electric dipole layer oscillating at the sum of the applied beams’ frequencies. In this approximation, the SFG signal is determined by a second-order electric dipole susceptibility [24, 25] and model calculations allow for the determination of the orientation of chemical bonds [26, 27, 28], the effective interfacial dielectric constant [29, 30, 31], the position of the dipole layer [29, 30], the surface potential [32, 33, 34, 35], and water structure around macromolecules [36, 37] from experiments. However, this dipole-layer picture neglects higher-order multipole contributions present in the SFG signal. Those multipole contributions to SFG spectra are proportional to higher-order response functions that are non-zero even in isotropic bulk media [38, 39, 40, 41, 42] and cannot be easily experimentally separated from the measured signal. To determine the structure of an interface from SFG spectra, one needs to identify the electric dipole component, which serves as a fingerprint of the interface structure. Thus, accurate theoretical predictions of higher-order multipole contributions are essential. Off-resonant multipole contributions, measured in second-harmonic generation, were predicted in previous works [43, 44]. Resonant multipole SFG contributions where estimated using normal-mode calculations [45] and found to be significant for the water bending band [46]. However, normal-mode methods involve rather drastic approximations, including locality assumptions, and cannot predict the spectral line shape. Experimentally, the multipolar origin of the water bending band has been investigated with SFG at surfactant monolayers [47, 48, 49] and combined SFG/DFG measurements [50], yielding conflicting results. Many studies attribute the bending band to the electric dipole component of the SFG signal [51, 52, 53, 47, 54, 55], in contrast to our findings.
Figure 1: SFG spectra and decomposition into multipole components. (a) & (b): Sketch of the second-order molecular multipole moments that give rise to SFG spectra. Here, $\mu_{i}^{(2)}$, $Q_{ij}^{(2)}$ and $m_{i}^{(2)}$ denote the induced molecular second-order electric dipole, electric quadrupole, and magnetic dipole moments, respectively. The contributions to $\tilde{S}^{(2)}_{yyz}$ are sketched in (a) and to $\tilde{S}^{(2)}_{zzz}$ in (b). The blue region represents water and the white region air. (c): Comparison between predicted and experimental SFG spectra $\tilde{S}^{(2)}_{yyz}$ in the bending frequency region. The SFG spectrum is decomposed into the pure electric dipole ($D$), the electric dipole - electric quadrupole cross (DQ), the electric quadrupole (Q), and the magnetic dipole contribution (M) in (e). (d) & (f): The same analysis for $\tilde{S}^{(2)}_{zzz}$. Results for the OH-stretch frequency region are shown in (g)-(j). Experimental data is taken from Fellows et al. [50], Yu et al. [30] and Chiang et al. [31]. To align with the experimental data, we red-shift our spectra by $28\text{\,}\mathrm{c}\mathrm{m}^{-1}$ in (c)-(f), and by $166\text{\,}\mathrm{c}\mathrm{m}^{-1}$ in (g)-(j). Grey dashed boxes indicate the regions shown in the insets.
Our method, based on time-dependent perturbation theory [56, 57], enables the quantitative prediction of multipolar SFG spectral contributions and constitutes a crucial step toward an accurate determination of the interfacial structure. We compute all relevant multipole contributions, including the magnetic dipole contributions, which were not considered in the past but are found to be sizable and of similar intensity as the electric dipole contribution in the bending band. Following the work of Guyot-Sionnest and Shen [58], we formulate the multipolar second-order electric current density as a response to electric displacement (D) fields perpendicular to the interface and electric (E) fields parallel to the interface, both of which are spatially constant on the relevant length scale. In doing so, we eliminate the need to make ad-hoc assumptions on the interface structure. We observe quantitative agreement between simulated and experimental SFG spectra, with significant multipole contributions. Recent experiments investigated the interfacial dielectric constant [31] and the spatial distribution of anisotropically oriented molecules [29, 59, 21]. Our work reveals that interfacial water exhibits an $8\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ thick biaxial triple-layer structure, as follows from the second-order electric-dipole susceptibility. This susceptibility is dominated by the biaxial orientational ordering perpendicular to the water molecular dipole axis in the bending region.
II Results
Figure 2: Spatial analysis of the bending band. The mass density profile is shown in (a), the local field factors $f^{\mathrm{VIS}}_{i}(z)$ defined in Equation (1) in (b). Lorentz theory predictions are denoted by dashed horizontal lines: black for vacuum, red and blue for bulk water. For illustrative purposes, a snapshot from the simulation is shown in the background. (c): Comparison of the spatially resolved integrals over the bending band, as defined in Equation (2), between $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ and the prediction based solely on molecular orientation, $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$, defined in Equation (33). The electric dipole second-order susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$, defined in Equation (27), is shown at selected positions in (d). These positions are marked in (f), where the corresponding two-dimensional profile is presented. The second-order response profile of the SFG signal $\tilde{s}^{(2)}_{yyz}(z)$ is presented in (e) and $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ in (g). The spectra are red-shifted by $28\text{\,}\mathrm{c}\mathrm{m}^{-1}$.
II.1 Quantitative Comparison of SFG Spectra with Experiments
In Figure 1, we present SFG spectra $\tilde{S}^{(2)}_{ijk}$ for the polarization combinations yyzyyz and zzzzzz, as defined in Equations (7) and (11). Here, the indices ijkijk denote the polarizations of the second-order electric current density, which produces the observed radiation, the VIS field, and the IR field, respectively. The theoretical framework for predicting multipolar SFG spectra is described in detail in Supplemental Material (SM) Sections I–VI [60]. We observe quantitative agreement between our predictions (red lines) and experiments (black lines) in the bending and stretching bands in Figure 1 (c), (g), (h). We analyze the SFG spectrum by dissecting it into its multipole contributions $\tilde{S}^{(2)}_{ijk}=\tilde{S}^{(2,\mathrm{DD})}_{ijk}+\tilde{S}^{(2,\mathrm{DQ})}_{ijk}+\tilde{S}^{(2,\mathrm{Q})}_{ijk}+\tilde{S}^{(2,\mathrm{M})}_{ijk}$. This decomposition is described in Appendix B. Here, $\tilde{S}^{(2,\mathrm{DD})}_{ijk}$ and $\tilde{S}^{(2,\mathrm{DQ})}_{ijk}$ are the pure electric dipole contribution and the electric dipole - electric quadrupole cross contribution, respectively. Importantly $\tilde{S}^{(2,\mathrm{DD})}_{ijk}$ describes the interfacial structure, while $\tilde{S}^{(2,\mathrm{DQ})}_{ijk}$ is created by the linear response of electric dipoles to second-order electric quadrupoles; $\tilde{S}^{(2,\mathrm{Q})}_{ijk}$ and $\tilde{S}^{(2,\mathrm{M})}_{ijk}$ are the electric quadrupole and magnetic dipole contributions, which are in contrast bulk contributions, independent of the interfacial structure. The sum of these contributions $\tilde{S}^{(2,\mathrm{Q})}_{ijk}+\tilde{S}^{(2,\mathrm{M})}_{ijk}$ is called the interfacial quadrupole bulk contribution (IQB) in the literature [41, 42]. To give some intuitive understanding of the physical mechanism producing the different multipolar SFG contributions we schematically illustrate the second-order molecular multipoles of the valence electron charge cloud in Figure 1 (a) and (b). Figure 1 (a) shows the contributions relevant for a y-polarized and (b) for a z-polarized SFG field. The second-order molecular electric dipole moment $\mu_{i}^{(2)}$ is characterized by oscillations of the displacement of the center of the electronic charge distribution and is zero in isotropic media. In contrast, the second-order electric quadrupole moment, characterized by oscillations of the charge distribution width, and the molecular magnetic dipole moment, characterized by oscillations in its angular momentum, as illustrated in Figure 1 (a) and (b), are nonzero in bulk. The second-order electric quadrupole and magnetic dipole moments consist of oppositely directed currents and contribute to the second-order current density via the gradients of their spatial distributions. The bending band SFG spectra in Figure 1 (c)-(f) are dominated by the electric quadrupole contribution (purple), which consists of a positive broad band. The pure electric dipole contribution consists of a negative-positive double peak (Fig. 1 (e) & (f)). The observed lineshape of $\tilde{S}^{(2,\mathrm{DD})}_{yyz}$ qualitatively aligns with previous simulation studies [51, 52, 54], which did not account for higher-order multipole contributions. The magnetic dipole contribution appears only in the yyzyyz spectrum and exhibits an intensity similar to the pure electric dipole contribution (Fig. 1 (e)). The OH-stretch bands are presented in Figure 1 (g)-(j). The pure electric dipole contribution in the stretching region is characterized by a negative–positive double peak, similar to the bending band, as can be seen by comparing the cyan lines in Figure 1 (e) & (i). The difference is that the frequency splitting between the positive and negative components is much larger compared to the spectral linewidth in the OH-stretch region, thus less cancellation takes place and the resulting electric dipole contribution is significantly stronger. The signal arising from the free OH stretch vibrations at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$ is largely determined by the pure electric dipole contribution (Fig. 1 (i) & (j)), while significant electric quadrupole contributions to the SFG spectra arise from the more slowly oscillating OH bonds between $3100\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and $3650\text{\,}\mathrm{c}\mathrm{m}^{-1}$. In $\tilde{S}^{(2)}_{yyz}$, the electric quadrupole produces most of the shoulder at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$ (Fig. 1 (i)). In all spectra presented, $\tilde{S}^{(2,\mathrm{DQ})}_{ijk}$ is small, and $\tilde{S}^{(2,\mathrm{DQ})}_{yyz}$ is mainly positive, while $\tilde{S}^{(2,\mathrm{DQ})}_{zzz}$ is mainly negative. All spectra have been red-shifted to ensure alignment between theory and experiment. This is reasonable since it has been demonstrated previously that the MB-Pol water model used by us reproduces the experimental vibrational frequencies when nuclear quantum effects are considered [61, 62, 63], while for the linear absorption spectra, it has been demonstrated that nuclear quantum effects do not alter the line shape and absolute intensities much [64, 62].
II.2 Spatially Resolved Second-Order Response Profile of the Bending Band
Figure 2 presents the position-dependent second-order response profile of the bending band, with $z$ given relative to the Gibbs dividing surface $z_{\mathrm{GDS}}$ [65]. Figure 2 (a) presents the mass density profile with a snapshot of the simulation box in the background. The density transitions from $1\text{\,}\mathrm{g}\mathrm{/}\mathrm{c}\mathrm{m}^{3}$ in bulk to zero over a range of about $5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$, centered around $z_{\mathrm{GDS}}$.
Figure 3: The orientation distribution function $\rho_{\mathrm{ORI}}(\theta,\psi)$ is shown for the same positions relative to $z_{\mathrm{GDS}}$ as in Figure 2 in (a)-(c). Three orientational species are distinguished: planar, pointing in, and pointing out, as illustrated at the top of panel (d). These orientations are marked with dots and contours in (a)-(c). Additionally, a snapshot of the simulation box viewed from the air is provided in (d), where molecules are color-coded according to their orientational species. (e) displays the fraction of each orientational species as a function of $z$. The molecular hyperpolarizability tensor $\tilde{\beta}_{yyz}(\theta,\psi)$, defined in Equation (3), is shown as a function of the molecular orientation in (f). The color indicates the intensity at bending-band frequencies, obtained by integrating the imaginary part $\tilde{\beta}^{\prime\prime}_{yyz}(\theta,\psi)$ using the same integration boundaries as in Equation (2). This intensity is decomposed into its uniaxial and biaxial components, defined in Equations (5) and (6) and shown in (g) and (h), respectively. The position-resolved profile of the bending band is decomposed into uniaxial and biaxial contributions in (i). The frequency-dependent coefficients $\tilde{\beta}^{lm}_{yyz}$ of $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}$ defined in Equation (4)-(6) are presented in (j).
We present the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$, which connects the average local electric E-field at the molecular centers to the resulting second-order electric dipole current density in Figure 2 (d) & (f). The superscript DL stands for dipole and local to distinguish it from regular susceptibilities, which are defined as the response of the macroscopic polarization density to macroscopic electric E-fields [66]. $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ enables linking the macroscopic second-order electric dipole response to molecular orientation [67, 24, 25]. In defining $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ via Equation (27), we account for the strength of the local E-field by dividing the second-order response of the electric dipole density $\tilde{s}^{(2,\mathrm{DD})}_{ijk}(z)$, by the local field factors, i.e. the ratios between the amplitudes of the local E-fields $\mathcal{E}^{\mathrm{L},\alpha}_{i}(z)$ acting on the molecular centers and the amplitudes of the external fields $\mathcal{F}^{\alpha}_{i}$, which correspond to D-fields for $i=z$ and to E-fields for $i=x$ or $y$. Here, $\alpha$ specifies the frequency of the three fields involved (SFG, VIS, IR). The relationship between $\tilde{s}^{(2,\mathrm{DD})}_{ijk}(z)$ and $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ is given in Equation (28). The spatial integral over the second-order response profile $\tilde{s}^{(2,\mathrm{DD})}_{ijk}(z)$ in Equation (18) determines the contribution of the pure electric dipole to the experimentally measurable SFG spectrum $\tilde{S}^{(2,\mathrm{DD})}_{ijk}$ presented in Figure 1 (c). The local field factors are defined by
$f^{\alpha}_{i}(z)=\frac{\mathcal{E}^{\mathrm{L,\alpha}}_{i}(z)}{\mathcal{F}^{\alpha}_{i}(z)}\,.$ (1)
The local field factor in the optical frequency range $f_{i}^{\mathrm{VIS}}(z)$ is presented in Figure 2 (b). Here, $f_{x}^{\mathrm{VIS}}(z)$ exhibits a line shape similar to the mass density profile. In bulk, we recover values dictated by the Lorentz field approximation [67, 25] $f^{\mathrm{VIS}}_{x}(-\infty)\approx(2+\varepsilon^{\mathrm{VIS}})/3$ and $f^{\mathrm{VIS}}_{z}(-\infty)\approx\left(2+\varepsilon^{\mathrm{VIS}}\right)/(3\varepsilon^{\mathrm{VIS}})$, as indicated by the horizontal dashed lines. Here, $\varepsilon^{\mathrm{VIS}}=1.77$ represents the optical dielectric constant of bulk water, extracted from the plateau value of the dielectric profile presented in Figure 5 (b). These findings are qualitatively consistent with the prediction of $f^{\mathrm{VIS}}_{i}(z)$ by Shiratori and Morita [68]. We compare the total second-order response function $\tilde{s}^{(2)}_{yyz}(z)$ defined in Equation (7) with the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ defined in Equation (27), as well as its prediction based solely on molecular orientation, $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$, in Figure 2 (e)-(g). It is evident that all profiles exhibit a positive contribution for values of $z-z_{\mathrm{GDS}}\geq 0$ and a negative contribution centered around $z-z_{\mathrm{GDS}}=$-2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$. We calculate $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ by assigning the molecular hyperpolarizability tensor to each water molecule in its molecular frame, as defined in Equation (33), where the molecular hyperpolarizability tensor is extracted from a simulation of bulk water as described in SM Section VII [60]. By this, the only interface-specific input to $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ is the molecular orientation distribution, allowing us to test whether $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ is a marker of the interfacial orientation anisotropy. To compare the amplitudes of $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ and $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ we integrate over the bending band according to
$\bar{\chi}^{(2,\mathrm{DL})}_{ijk}(z)=c_{0}\int\limits_{\nu_{1}}^{\nu_{2}}\mathrm{d}\nu\,\tilde{\chi}^{{(2,\mathrm{DL}})\prime\prime}_{ijk}(z)$ (2)
from $\nu_{1}=$1507\text{\,}\mathrm{c}\mathrm{m}^{-1}$\$ to $\nu_{2}=$1772\text{\,}\mathrm{c}\mathrm{m}^{-1}$\$. Here, $c_{0}$ is the speed of light in vacuum. We find quantitative agreement between $\bar{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ and $\bar{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ in Figure 2 (c), demonstrating that $\bar{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ indeed primarily reflects interfacial orientation within the bending frequency range. The characteristic triple-layer structure of $\bar{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ will be discusses further below. By comparing $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$ (Fig. 2 (d) & (f)) with $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ (Fig. 2 (g)), it becomes clear that $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ does not reproduce the $z$-dependent frequency shift, present in $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}(z)$, which is thus due to the $z$-dependent anisotropic coordination of interfacial water.
II.3 Spectral Signature of Biaxial Interfacial Water Ordering
After demonstrating that the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{yyz}$ serves as a quantitative probe of the interfacial orientation of molecules, we now analyze the interfacial orientation distribution. The orientation distribution function $\rho_{\mathrm{ORI}}(\theta,\psi)$, defined in Equation (34), is presented at three positions relative to $z_{\mathrm{GDS}}$ in Figure 3 (a)-(c). This function depends on the two Euler angles, $\theta$ and $\psi$. The angle $\theta$ describes the tilt of the molecular dipole axis relative to the surface normal: for $\theta=0$, the oxygen atom is oriented towards bulk water, and for $\theta=180^{\circ}$ towards air. The angle $\psi$ describes the molecular orientation around the dipole axis. The interfacial water structure is significantly biaxial, as revealed by the pronounced $\psi$-dependence of $\rho_{\mathrm{ORI}}(\theta,\psi)$. We distinguish three orientational species: the planar-oriented species (orange), where both OH-bond vectors lie in the interfacial plane, the inward-oriented species (magenta), where one OH-bond is orthogonal to the interfacial plane and pointing inward, and the outward-oriented species (cyan), where one OH-bond is orthogonal to the interfacial plane and pointing outward. These orientations are sketched on top of Figure 3 (d) and the corresponding Euler angles are marked with color-coded dots and contours (marking water molecules with angular deviation of $36^{\circ}$ from the idealized orientations) in the plots of $\rho_{\mathrm{ORI}}(\theta,\psi)$ in Figure 3 (a)-(c). A snapshot of the air-water interface viewed from air is presented in Figure 3 (d); here, the orientational species are color-coded. As observed in Figure 3 (a), at $2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ relative to $z_{\mathrm{GDS}}$, the inward-oriented species dominates; at $z=z_{\mathrm{GDS}}$, the planar species prevails (Fig. 3 (b)), and finally, close to the vapor region at $1.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$, the outward-oriented species becomes dominant (Fig. 3 (c)). We quantify this observation via the molecular number fraction profiles, which we present in Figure 3 (e). Here, we see that water is isotropic $5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ below $z_{\mathrm{GDS}}$, as the fraction of all orientational species is approximately $20\text{\,}\%$. The inward-oriented fraction peaks at $2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$, at the location of the negative peak of $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}(z)$ in Figure 2 (f). The planar-oriented and outward-oriented species peak at $z_{\mathrm{GDS}}$ and near the vapor region at $2.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$, respectively. The uniaxial approximation neglects the biaxial water ordering, which corresponds to averaging the orientation distribution function over the angle $\psi$. To reveal the spectral effects of water biaxial ordering, we define the orientation and position dependent molecular hyperpolarizability $\tilde{\beta}_{ijk}(z,\theta,\psi)$ by
$\varepsilon_{0}^{-1}\mu_{i}^{(2)}(t)=e^{-i\omega^{\mathrm{SFG}}t}\tilde{\beta}_{ijk}(z,\theta,\psi)\mathcal{E}^{\mathrm{L,VIS}}_{j}(z)\mathcal{E}^{\mathrm{L,IR}}_{k}(z)\\ +c.c.\,$, (3)
where $\mu_{i}^{(2)}(t)$ is the second-order molecular electric dipole moment, $\mathcal{E}^{\mathrm{L,VIS}}_{j}(z)$ and $\mathcal{E}^{\mathrm{L,IR}}_{k}(z)$ are the amplitudes of the local E-fields acting on the molecular center at the position $z$, and $\theta$ and $\psi$ are the Euler angles, specifying the molecular orientation. Note that $\mu_{i}^{(2)}(t)$ is a source dipole that induces an additional linear response, as explained in SM Section IV B [60]. We extract $\tilde{\beta}_{yyz}(\theta,\psi)$, denoting the bulk water hyperpolarizability, from separate trajectories of bulk water, as described in SM Section VII [60]. In Figure 3 (f) we present the integral of the imaginary part $\tilde{\beta}^{\prime\prime}_{yyz}(\theta,\psi)$ over the same integration boundaries as in Equation (2), to quantify the orientation-dependent hyperpolarizability in the bending band. We decompose $\tilde{\beta}_{yyz}(\theta,\psi)$ into uniaxial and biaxial components according to
$\tilde{\beta}_{ijk}(\theta,\psi)$ $=\tilde{\beta}^{\mathrm{UNI}}_{ijk}(\theta)+\tilde{\beta}^{\mathrm{BI}}_{ijk}(\theta,\psi)$ (4)
$\tilde{\beta}^{\mathrm{UNI}}_{ijk}(\theta)$ $=q_{10}(\theta)\tilde{\beta}^{10}_{ijk}+q_{30}(\theta)\tilde{\beta}^{30}_{ijk}$ (5)
$\tilde{\beta}^{\mathrm{BI}}_{ijk}(\theta,\psi)$ $=q_{32}(\theta,\psi)\tilde{\beta}^{32}_{ijk}$ (6)
shown in Figure 3 (g) & (h), respectively. Here, $q_{10}(\theta)=\cos{\theta}$, $q_{30}(\theta)=\frac{1}{2}\left[5\cos{{}^{3}\theta}-3\cos{\theta}\right]$ are the first and third Legendre polynomials specifying the dipole distribution and $q_{32}(\theta,\psi)=\cos{\theta}\sin{\theta}^{2}\left(\cos{{}^{2}\psi}-\sin{{}^{2}\psi}\right)$ accounts for the contribution due to the molecular biaxiality. Equations (4)–(6) express the rotation of the molecular hyperpolarizability into the laboratory frame in an exact way. The frequency-dependent imaginary parts of the coefficients $\tilde{\beta}^{10}_{yyz}$, $\tilde{\beta}^{30}_{yyz}$ and $\tilde{\beta}^{32}_{yyz}$ are presented in Figure 3 (j), where we observe that the biaxial contribution $\tilde{\beta}^{32}_{yyz}$ is significant. We observe that the negative area at $\theta>90^{\circ}$ in $\tilde{\beta}_{yyz}(\theta,\psi)$ in Figure 3 (f) does align well with the inward pointing orientation dominant at $2\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ in Figure 3 (a) and the positive area at $\theta<90^{\circ}$ in Figure 3 (f) does align with the outwards pointing orientation dominant at $1.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ in Figure 3 (c). We find that the uniaxial contribution in Figure 3 (g) for each orientational species cancels out, meaning that uniaxial water ordering does not lead by itself to an SFG signal in the bending region. In contrast, the biaxial component projects precisely out the difference between outwards- and inwards-oriented molecules (Fig. 3 (h)). We note that planar-oriented molecules are not detected by SFG spectroscopy because their mirror plane lies parallel to the interface, making them inversion symmetric. We dissect the integral over the bending band of $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$, already presented in Figure 2 (c), into biaxial and uniaxial contributions in Figure 3 (i), demonstrating that $\tilde{\chi}^{(2,\mathrm{ORI})}_{yyz}(z)$ is dominated by the biaxiality of the orientation distribution.
II.4 Spatially Resolved Second-Order Response Profile of the Stretch Band
Figure 4: Second-order response profile $\tilde{s}^{(2)}_{ijk}\left(z\right)$ of the OH-stretch band as defined in Equation (7) and the decomposition into its multipole components $\tilde{s}^{(2,\beta)}_{ijk}\left(z\right)$ as defined in Equation (17). A snapshot of the simulation box is shown in (a). The mass-density profiles with respect to $z_{\mathrm{GDS}}$ and $z_{\mathrm{WCS}}$ are presented in (b). Slices of $\tilde{s}^{(2)}_{ijk}\left(z\right)$ are shown for selected positions in (c) & (d) relative to $z_{\mathrm{GDS}}$ and in (e) & (f) relative to $z_{\mathrm{WCS}}$, colored vertical lines in (a), (b), (g)-(j) indicate the positions. The profiles $\tilde{s}^{(2)}_{ijk}\left(z\right)$ are shown relative to $z_{\mathrm{WCS}}$ in (g) & (h) and relative to $z_{\mathrm{GDS}}$ in (i) & (j). The profile relative to $z_{\mathrm{GDS}}$ is dissected into its molecular multipole components $\tilde{s}^{(2,\mathrm{D})}_{ijk}\left(z\right)$ and $\tilde{s}^{(2,\mathrm{Q})}_{ijk}\left(z\right)$ in (k) & (l) and (m) & (n), respectively. The spectra are red-shifted by $166\text{\,}\mathrm{c}\mathrm{m}^{-1}$.
We present the imaginary part of the second-order response profile $\tilde{s}^{(2)}_{ijk}(z)$ in the OH-stretch frequency region, defined in Equation (7), together with a decomposition into the electric dipole and electric quadrupole contribution according to Equation (17) in Figure 4. We map the waters center of mass positions in the calculation of $\tilde{s}^{(2)}_{ijk}(z)$ relative to the non-planar Willard-Chandler surface $z_{\mathrm{WCS}}$ [69] in Figure 4 (e)-(h) and relative to the planar Gibbs dividing surface $z_{\mathrm{GDS}}$ in Figure 4 (c) & (d) and (i)-(n). We show a snapshot of the simulation box in Figure 4 (a). In Figure 4 (b), we present the mass-density profiles relative to $z_{\mathrm{GDS}}$ and $z_{\mathrm{WCS}}$. In the latter presentation the density profile strongly oscillates, while in the laboratory frame these oscillations are washed out due to the intrinsic roughness of the air-water interface [70]. Slices of $\tilde{s}^{(2)}_{ijk}(z)$ at selected positions $z-z_{\mathrm{GDS}}$ and $z-z_{\mathrm{WCS}}$ (denoted by the colored lines in (b)) are presented in Figure 4 (c) & (d) and (e) & (f). Close to the vapor region at $z-z_{\mathrm{GDS/WCS}}=$0.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ (orange lines), almost only the free OH stretch peak is visible at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$. At $z-z_{\mathrm{GDS/WCS}}=$-0.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ (purple lines), $\tilde{s}^{(2)}_{ijk}(z)$ qualitatively agrees with the integrated spectra presented in Figure 1 (g) & (h). Finally, the free OH stretch contributions vanish near the bulk region at $z-z_{\mathrm{GDS/WCS}}=$-1.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ (turquoise lines), and we only see the shoulder at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and the negative component at $3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$. As shown in Figure 4 (h), $\tilde{s}^{(2)}_{zzz}(z)$ relative to $z_{\mathrm{WCS}}$ almost perfectly follows the oscillations of the mass density profile in Figure 4 (b). The second-order response profile in the laboratory frame $\tilde{s}^{(2)}_{ijk}(z)$ is presented in Figure 4 (c) & (d) and (i) & (j). Here, both the free OH stretch band at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and the negative contribution at $3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$ are located at approximately $z\approx z_{\mathrm{GDS}}$. However, the shoulder is created below $z_{\mathrm{GDS}}$, and the free-OH contributions reaches slightly more into the vapor region. The profiles of the second-order response of the electric dipole contribution $\tilde{s}^{(2,\mathrm{D})}_{ijk}(z)$ are presented in Figure 4 (k) & (l). These qualitatively agree with previously published depth-resolved SFG spectra [71, 72, 73, 74, 21, 75, 76]. The second-order electric dipole contributions in the stretching band (Fig. 4 (k) & (l)) and in the bending band (Fig. 2 (f)) undergo significant frequency shifts as a function of position $z$ in the interfacial layer. As shown elsewhere [77], when going from vapor to liquid, frequency shifts arise from competing effects of non-Markovian friction (which causes blue-shifting) and potential broadening (which causes red-shifting). The former dominates in the bending band. The electric quadrupole profile $\tilde{s}^{(2,\mathrm{Q})}_{ijk}(z)$ is presented in Figure 4 (m) & (n) and is dominated by a broad positive peak around $z\approx z_{\mathrm{GDS}}$. Additionally, we observe a negative contribution closer to the bulk region and a positive contribution closer to the vapor region at the frequency of the free OH vibrations. It is important to note that the electric quadrupole profile $\tilde{s}^{(2,\mathrm{Q})}_{ijk}(z)$ strongly depends on position. Consequently, the electric quadrupole contribution to the SFG spectra, $\tilde{S}^{(2,\mathrm{Q})}_{ijk}$ as defined in Equation (18), does report on interfacial water structure for a non-vanishing wave vector mismatch $\Delta k_{z}$, which is an interesting venue for future experimental investigation.
II.5 Spatially Resolved Linear Dielectric and Absorption Profile
Figure 5: Linear dielectric and absorption profiles at the air water interface. (a): Snapshot of the simulation. (b): The density profile is compared to the real part of the optical dielectric profile. (c) & (d): The linear absorption profiles $\omega\tilde{\varepsilon}^{\prime\prime}_{xx}(z)$ and $\omega\tilde{\varepsilon}^{\prime\prime}_{zz}(z)$ in the IR frequency range are shown as a 2D plot. (e)-(j): The imaginary part of the linear absorption profile is shown at the same positions relative to the Gibbs dividing surface as in Figure 4. The xxxx-component defined in Equation (25), the zzzz-component defined in Equation (24), and the isotropic component defined in (26) are presented in (e), (f), and (g), respectively. Additionally, the predicted linear absorption spectrum of bulk water and the experimental spectrum [78] are shown in (g). (h)-(j): Results in the stretch region. The bending band in (e)-(g) is redshifted by $28\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and the stretching band in (h)-(j) is red-shifted by $166\text{\,}\mathrm{c}\mathrm{m}^{-1}$.
As mentioned before, the $z$-independent external field amplitude $\mathcal{F}^{\alpha}_{z}$ corresponds to the D-field, while $\mathcal{F}^{\alpha}_{x/y}$ corresponds to the E-field. The relationship between D- and E-fields is determined by the dielectric profile [79, 80, 81, 82, 83, 84, 85]. The real part of the dielectric profile $\tilde{\varepsilon}_{xx}^{\mathrm{VIS}}(z)$ (red) and $\tilde{\varepsilon}_{zz}^{\mathrm{VIS}}(z)$ (blue) at optical frequencies is presented in Figure 5 (b) and has an almost identical shape to the mass density profile (black). Thus, in contrast to the static or IR case [80, 86], the anisotropy of the tensorial interfacial response for optical frequencies is rather small, as demonstrated in the inset. The bulk plateau value of $1.77$ agrees well with the experimental value of $1.78$ at room temperature under atmospheric pressure [87]. We present the imaginary part of the position and frequency-dependent dielectric profiles, $\tilde{\varepsilon}^{\mathrm{IR}}_{xx}(z)$ and $\tilde{\varepsilon}^{\mathrm{IR}}_{zz}(z)$ extracted from molecular dynamics simulations using Equations (24) and (25) in Figure 5 (c) & (d). We observe that the dielectric profile $\tilde{\varepsilon}^{\mathrm{IR}}_{ij}(z)$ varies drastically in the interface region and is significantly anisotropic. We see that the change in the parallel component of the dielectric profile $\tilde{\varepsilon}^{\mathrm{IR}}_{xx}\left(z\right)$ (Fig. 5 (c) & (e) & (h)) is mainly characterized by a frequency-independent decrease of intensity across the interface, without significant changes in the line shape. The transition of the perpendicular component $\tilde{\varepsilon}_{zz}(z)$ (Fig. 5 (d) & (f) & (i)) is more complex. The OH-stretch band, which is in bulk centered around $3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$, undergoes a blue shift from the bulk to vapor until only the free OH-stretch part at $3700\text{\,}\mathrm{c}\mathrm{m}^{-1}$ survives. The libration band [88, 89], between $250\text{\,}\mathrm{c}\mathrm{m}^{-1}$ and $1000\text{\,}\mathrm{c}\mathrm{m}^{-1}$, is more pronounced in $\tilde{\varepsilon}^{\mathrm{IR}}_{zz}(z)$ than in $\tilde{\varepsilon}^{\mathrm{IR}}_{xx}(z)$, as can be seen in Figures 5 (e) & (f). This can be explained by the anisotropic orientation of the molecular dipole moments, and likely results from a higher concentration of planar-oriented molecules near the interface (Fig. 3 (e)). We observe that the interfacial and the bulk spectra are indistinguishable at $z-z_{\mathrm{GDS}}=$-5.5\text{\,}\mathrm{\SIUnitSymbolAngstrom}$\$ in Figure 5 (g) & (j). Hence, the second-order response in Figures 2 & 4, as well as the linear response in Figure 5, and the orientational anisotropy in Figure 3, are bulk-like at this depth. We compare with experiments [78] (black) in Figure 5 (g) & (j), where we observe that the differences between simulation and experiments are rather small in the wavenumber range presented in Figure 5 (g). Especially, the difference between the experimental bending peak and simulation is small. In the OH-stretch region in Figure 5 (j), we observe good agreement at high-frequencies ($3500\text{\,}\mathrm{c}\mathrm{m}^{-1}$\$) and an underestimation of the experimental IR intensity at lower frequencies.
III Conclusion
Accounting for higher-order multipole contributions is essential for interpreting SFG spectra and leads to quantitative agreement between simulations and experimental data. In the bending band, the electric dipole contribution is almost negligible due to the small frequency shift between mutually canceling positive and negative SFG signals from water at different depths, and has an amplitude comparable to the previously neglected magnetic dipole contribution. As a result, the total SFG signal is dominated by the electric quadrupole contribution. The shoulder appearing at $3600\text{\,}\mathrm{c}\mathrm{m}^{-1}$ in the stretch band has been attributed to orientational anisotropy [90, 91], or to a combination mode involving intermolecular coupling [92], as summarized in the review by Tang et al. [93]. Our results demonstrate that the shoulder primarily originates from electric quadrupole contributions, which do not reflect interfacial structure but rather depend on bulk properties for vanishing wave vector mismatch $\Delta k_{z}=0$. After subtracting higher-order multipoles, the anisotropic orientation of interfacial molecules determines the SFG signal which reveals strong biaxial water ordering at the air-water interface. This biaxial structure is composed of a layering of essentially three orientational species, namely an inwards-oriented, a planar-oriented, and an outwards-oriented species.
In conclustion, since SFG spectra are significantly influenced by bulk-specific multipole effects, it is crucial to account for and subtract these contributions to interpret experimental spectra in terms of interfacial orientations. This can be achieved by subtracting a reference spectrum with the same bulk medium [50], or by using theoretical predictions, as those provided in this work.
IV Methods
All simulations are done in the NVT ensemble at room temperature $T=$298\text{\,}\mathrm{K}$\$ with the MB-Pol force field [94, 95, 96]. For starting configurations, 94 equally spaced initial configurations are extracted from a $18.8\text{\,}\mathrm{n}\mathrm{s}$ long simulation using the SPC/E force field [97] and the GROMACS molecular dynamics software [98]. The SPC/E simulations are done with a time step of $2\text{\,}\mathrm{f}\mathrm{s}$ and a velocity rescaling thermostat [99], with a relaxation time of $1\text{\,}\mathrm{p}\mathrm{s}$. Each of the initial configurations is equilibrated with the MB-Pol force field for $20\text{\,}\mathrm{p}\mathrm{s}$, and afterward, production runs with an average runtime of $0.86\text{\,}\mathrm{n}\mathrm{s}$ are executed using the LAMMPS software [100]. Again, a velocity-rescaling thermostat is employed. However, for MB-Pol simulations, the relaxation time is set to $5\text{\,}\mathrm{p}\mathrm{s}$, and the time step is $0.2\text{\,}\mathrm{f}\mathrm{s}$. The box size is $2\text{\,}\mathrm{n}\mathrm{m}$ in the x and y dimensions and $6\text{\,}\mathrm{n}\mathrm{m}$ in the z dimension. This size is chosen to ensure that the box contains more than twice as much air as water, since the water slab spans approximately $2.6\text{\,}\mathrm{n}\mathrm{m}$. The box is filled with $352$ water molecules and the two interfaces lie in the $xy$ plane. The total dipole moment of the simulation box is computed with the modified TTMF-4 model, which describes the electrostatics of the MB-Pol potential [64, 94]. The same model is used to predict the linear absorption profile, where we compute polarization density trajectories from monopole and dipole density trajectories. Molecular polarizabilities are computed with single-molecule quantum chemistry calculations using the software Gaussian 16 [101]. The theoretical level is CCSD(T)/aug-cc-pVTZ and B3LYP/aug-cc-pVTZ in the parameterization of $\alpha^{n,\mathrm{DD}}_{ij}(t)$ and $\alpha^{n,\mathrm{QD}}_{ijk}(t)$, appearing in Equations (19) and (20), respectively. Electrostatic interactions between the induced multipoles are computed as described in SM Sections V B and VIII [60], involving the Ewald summation algorithm included in OpenMM [102]. The simulation of bulk water is done in a cubic box with a length of $2.0727\text{\,}\mathrm{n}\mathrm{m}$ consisting of 297 molecules. The magnetic dipole contribution is extracted from a different trajectory, because it requires a much smaller write-out frequency. The initial configurations for the trajectories used to extract the bulk dielectric constant, molecular hyperpolarizability, and magnetic dipole contribution are taken from a trajectory generated with the SPC/E model after a simulation time of $20\text{\,}\mathrm{n}\mathrm{s}$ and $40\text{\,}\mathrm{n}\mathrm{s}$, respectively. The first $10\text{\,}\mathrm{p}\mathrm{s}$ of the MB-Pol trajectory is discarded, and a trajectory of length of $400\text{\,}\mathrm{p}\mathrm{s}$ is used to calculate the bulk dielectric constant and the molecular hyperpolarizability, and a trajectory of length of $300\text{\,}\mathrm{p}\mathrm{s}$ to calculate the second-order magnetic dipole susceptibility. The simulation parameters of the bulk simulations are the same as for the interface. We use the recommended settings of the MB-Pol potential from the GitHub repository "MBX - Version 0.7" from the Paesani group [103]. The write-out frequency is $3.2\text{\,}\mathrm{f}\mathrm{s}$ for all spectra except $0.4\text{\,}\mathrm{f}\mathrm{s}$ for the magnetic dipole contribution. In all simulations, periodic boundary conditions are imposed. However, we correct for the external field modification arising from the periodic replicas as described in SM Section IX [60], previously applied for the calculation of dielectric profiles [79, 68, 85]. The smoothing procedure is described in SM Section X [60].
V Acknowledgements
We acknowledge financial support from the Deutsche Forschungsgemeinschaft (DFG) within the grant CRC 1349 and HPC resources on CURTA [104], TRON, and Sheldon from the FU Berlin for data acquisition.
Appendix A Prediction of Locally Resolved SFG Spectra
We consider a planar interface between two media. The system is homogeneous in the $xy$ plane but inhomogeneous along the $z$ dimension. The system is illuminated by two monochromatic light sources with frequencies $\omega^{\mathrm{VIS}}$ and $\omega^{\mathrm{IR}}$ that induce an electric current density $j_{i}^{(2)}(z,t)$, oscillating with the sum frequency $\omega^{\mathrm{SFG}}=\omega^{\mathrm{IR}}+\omega^{\mathrm{VIS}}$. Here $\omega^{\mathrm{IR}}$ is in the IR range and invokes equifrequent oscillation of the nuclei, while $\omega^{\mathrm{VIS}}$ is in the optical frequency range. The resulting second-order electric current density is determined by the second-order response of a time-dependent perturbation expansion [57]
$\varepsilon_{0}^{-1}j_{i}^{(2)}(z,t)=-i\omega^{\mathrm{SFG}}e^{-i\omega^{\mathrm{SFG}}t}\tilde{s}^{(2)}_{ijk}\left(z\right)\mathcal{F}^{\mathrm{VIS}}_{j}\mathcal{F}^{\mathrm{IR}}_{k}+c.c.\,$, (7)
[…]
$\mathcal{F}^{\alpha}_{i}=\left(\delta_{ix}+\delta_{iy}\right)\mathcal{E}^{\alpha}_{i}+\delta_{iz}\varepsilon_{0}^{-1}\mathcal{D}^{\alpha}_{i}\,$, (8)
where $\mathcal{E}^{\alpha}_{i}$ and $\mathcal{D}^{\alpha}_{i}$ are the amplitudes of the E-fields and the D-fields
$E^{\mathrm{\alpha}}_{i}(z,t)$ $=\mathcal{E}^{\alpha}_{i}(z)e^{-i\omega^{\mathrm{\alpha}}t}+c.c.$ (9)
$D^{\mathrm{\alpha}}_{i}(z,t)$ $=\mathcal{D}^{\alpha}_{i}(z)e^{-i\omega^{\mathrm{\alpha}}t}+c.c.\,$, (10)
oscillating with the frequency $\omega^{\alpha}$, $\alpha\in\{\mathrm{SFG},\mathrm{VIS},\mathrm{IR}\}$ and $\delta_{ij}$ is the Kronecker delta. Hence, the symbol $\mathcal{F}^{\alpha}_{i}$ is a placeholder for the spatially constant amplitude of D-fields or E-fields on the relevant length scale, depending on the polarization. The external fields are given by $F^{\mathrm{\alpha}}_{i}(t)=\mathcal{F}^{\mathrm{\alpha}}_{i}e^{-i\omega^{\mathrm{\alpha}}t}+c.c.$}. By formulating the theory with respect to these fields, we avoid locality approximations. We note that $\tilde{s}^{(2)}_{ijk}(z)$ is distinct from the second-order susceptibility $\tilde{\chi}^{(2)}_{ijk}(z)$, which defines the response to electric E-fields. Here, we only consider the signal arising from the interface on the length scale at which the external fields are spatially constant. External field gradients lead to so-called bulk quadrupole contributions [41, 25, 42], which can be experimentally estimated using combined transmission-reflection SFG [105] or combined SFG/DFG techniques [21, 50] and are discussed in SM Section IV C [60]. The SFG spectrum follows by integration over $z$ as
$\tilde{S}^{(2)}_{ijk}$ $=\int\limits_{-\infty}^{\infty}\mathrm{d}z\,e^{-i\Delta k_{z}z}\tilde{s}^{(2)}_{ijk}(z)\,$, (11)
where $\Delta k_{z}$ is the wave vector mismatch defined in SM Section II [60]. We emphasize that Equation (11) holds whenever the external fields can be assumed to be constant in the region where the dielectric profile is inhomogeneous. This assumption is well satisfied in our case, because the shortest wavelength considered is $\lambda^{\mathrm{SFG}}\approx$600\text{\,}\mathrm{n}\mathrm{m}$\$}, and the region where the dielectric profile is inhomogeneous is only a few angstroms thick, as demonstrated in this work. Here we utilize the additional approximation $e^{-i\Delta k_{z}z}\tilde{s}^{(2)}_{ijk}(z)\approx\tilde{s}^{(2)}_{ijk}(z)$. At the air waiter interface, this is a good approximation, as a typical value for $\Delta k_{z}$ is $0.02\text{\,}\mathrm{n}\mathrm{m}^{-1}$, whereas the thickness of the air-water interface is about $8\text{\,}\mathrm{\SIUnitSymbolAngstrom}$ [21]. A derivation of Equations (7)-(11) is provided in SM Sections I and II [60].
Appendix B Molecular Multipole Contributions
The second-order electric current density in Equation (7) can be decomposed into molecular multipole contributions according to
$j^{(2)}_{i}(z,t)=j^{(2,\mathrm{D})}_{i}(z,t)+j^{(2,\mathrm{Q})}_{i}(z,t)+j^{(2,\mathrm{M})}_{i}(z,t)+$...\,, (12)
[…]
and
[…]
$\tilde{s}^{(2)}_{ijk}(z)$ $=\tilde{s}^{(2,\mathrm{D})}_{ijk}(z)+\tilde{s}^{(2,\mathrm{Q})}_{ijk}(z)+\tilde{s}^{(2,\mathrm{M})}_{ijk}(z)$ (15)
$\tilde{s}^{(2,\mathrm{D})}_{ijk}(z)$ $=\tilde{s}^{(2,\mathrm{DD})}_{ijk}(z)+\tilde{s}^{(2,\mathrm{DQ})}_{ijk}(z)\,$, (16)
where each contribution $\tilde{s}^{(2,\beta)}_{ijk}\left(z\right)$ is defined as
$\varepsilon_{0}^{-1}j_{i}^{(2,\beta)}(z,t)=-i\omega^{\mathrm{SFG}}e^{-i\omega^{\mathrm{SFG}}t}\tilde{s}^{(2,\beta)}_{ijk}\left(z\right)\mathcal{F}^{\mathrm{VIS}}_{j}\mathcal{F}^{\mathrm{IR}}_{k}+c.c.\,$, (17)
where $\beta\in\{\mathrm{DD},\mathrm{DQ},\mathrm{D},\mathrm{Q},\mathrm{M}\}$. Consequently, the spatially integrated multipole contributions to the SFG spectrum are given by
$\tilde{S}^{(2,\beta)}_{ijk}=\int\limits_{-\infty}^{\infty}\mathrm{d}z\,e^{-i\Delta k_{z}z}\tilde{s}^{(2,\beta)}_{ijk}\left(z\right)\,.$ (18)
While the total current $j^{(2)}_{i}(z,t)$ does not depend on the choice of the molecular origin, the individual contributions in Equation (12) do [108, 74]. To determine the optimal position of the molecular origin, we consider the SFG signal from an interface with isotropically oriented molecules $\tilde{S}^{(2,\mathrm{ISO})}_{ijk}$, which we call an isotropic interface. This isotropic interface is created by cutting bulk water at an arbitrary $z$ position into two halves. Consequently, boundary contributions created by the change in density at the interface are present, but the molecules’ orientational distribution is inversion symmetric. A similar interface was created to predict the Bethe potential [109]. We test three molecular origins for the calculation of multipoles and find that if we choose the molecular center of mass as the molecular origin, $\tilde{S}^{(2,\mathrm{ISO})}_{ijk}\approx\tilde{S}^{(2,\mathrm{Q})}_{ijk}+\tilde{S}^{(2,\mathrm{M})}_{ijk}$} does hold in good approximation. Thus, choosing the molecular center of mass as the expansion center ensures that the electric dipole contribution is solely due to molecular orientation, as discussed in detail in SM Section VI [60]. In contrast, choosing the molecular center different from the center of mass introduces significant boundary contributions in the electric dipole contribution $\tilde{S}^{(2,\mathrm{D})}_{ijk}$. Moreover, the center of mass maximizes the decoupling of molecular translations, vibrations, and rotations [110, 111, 112]. The theory of our multipole decomposition is described in SM Sections III and IV [60]. There, we also derive the constitutive relations, including nonlinear multipolar source terms, using the Lorentz field approximation in planar geometry, based on works by Mizrahi and Sipe [113] and Hirano and Morita [42]. These equations are solely needed to aid the interpretation of experimental spectra.
Appendix C Fluctuation-Dissipation Relations within the Off-Resonant Approximation
We assume that the VIS field interacts off-resonantly with the system, meaning it does polarize the molecules but does neither excite higher electronic levels nor influence the nuclei trajectories. This approximation is valid since the VIS field oscillates too rapidly to influence the motion of the nuclei yet slowly enough to act adiabatically on the distribution of the electrons. In this limit the SFG signal arises from a first-order perturbation expansion with respect to the external field from the IR light source [114]. The molecular multipoles are determined by the equations
$\mu^{n}_{i}(t)$ $=\alpha^{n,\mathrm{DD}}_{ij}(t)f^{n}_{jk}(t)F^{\mathrm{VIS}}_{k}(t)$ (19)
$Q^{n}_{ij}(t)$ $=\alpha^{n,\mathrm{QD}}_{ijk}(t)f^{n}_{kl}(t)F^{\mathrm{VIS}}_{l}(t)\,$, (20)
where $\mu^{n}_{i}(t)$ and $Q^{n}_{ij}(t)$ are the induced electric dipole and electric quadrupole moments of the nth-molecule, and $\alpha^{n,\mathrm{DD}}_{ij}(t)$ and $\alpha^{n,\mathrm{QD}}_{ijk}(t)$ are the electric dipole and the electric quadrupole polarizabilities of the $n^{\mathrm{th}}$ molecule. The local field factor $f_{ij}^{n}(t)$ transforms an external field $F_{i}(t)$ into the local E-field acting on the nth molecule, $E^{n}_{i}(t)=f_{ij}^{n}(t)F_{j}(t)$}, and is determined in a self-consistent manner such that the multipoles induced by $E^{n}_{i}(t)$ produce the field $E^{n}_{i}(t)-F_{i}(t)$. It is demonstrated in SM Section V B [60] that it is incorrect to solve separate self-consistent field equations for the VIS and the SFG field in the time domain, as proposed previously [114, 115, 116]. Rather, if a timescale-separation approximation is applied, Equations (19) and (20) each split into components at SFG and VIS frequencies, as described in SM Section V B [60]. The polarization contributions of the second-order response profiles, defined in Equation (17), are determined by a first-order perturbation expansion, for which we introduce the fluctuation-dissipation relations
$s^{(2,\mathrm{D})}_{ijk}(z,t)$ $=\frac{-\Theta(t)}{Ak_{B}T\varepsilon_{0}}\frac{\partial}{\partial t}\sum\limits_{n}^{N_{\mathrm{mol}}}\left\langle\alpha^{n,\mathrm{DD}}_{il}(t)f^{n}_{lj}(t)\delta\left[z-z^{n}(t)\right]P_{k}(0)\right\rangle$ (21)
$s^{(2,\mathrm{Q})}_{ijk}(z,t)$ $=\frac{\Theta(t)}{Ak_{B}T\varepsilon_{0}}\frac{\partial}{\partial z}\frac{\partial}{\partial t}\sum\limits_{n}^{N_{\mathrm{mol}}}\left\langle\alpha^{n,\mathrm{QD}}_{izl}(t)f^{n}_{lj}(t)\delta\left[z-z^{n}(t)\right]P_{k}(0)\right\rangle,$ (22)
where $\langle$$\rangle$ denotes ensemble averaging, $T$ is the temperature, $k_{B}$ the Boltzmann constant, $\theta(x)$ denotes the Heaviside function and the z-position of the $n^{\mathrm{th}}$ molecule is denoted as $z^{n}$. These expressions omit the off-resonant hyperpolarizability, which can be included as shown in SM Section V [60], but does not contribute to the imaginary part of SFG spectra. Equations (21) and (22) follow from time-dependent perturbation theory using the perturbation Hamiltonian $H_{\mathrm{int}}(t)=-P_{i}F^{\mathrm{IR}}_{i}(t)$}, where $P_{i}$ is the total system’s dipole moment. Hirano and Morita [74] suggest the perturbation expansion using the modified perturbation Hamiltonian $H_{\mathrm{int}}^{+}(t)=-P^{+}_{i}F^{\mathrm{IR}}_{i}(t)$}, where $P^{+}_{i}(t)$ is the dipole moment of the upper half of the simulation box. It is demonstrated in SM Section V D [60] that this introduces a spurious contribution from the interface between the two halves. We note that while we treat electric dipole and electric quadrupole contributions in a precise manner, we approximate magnetic dipole contributions by considering only the leading order term of an expansion in terms of electric multipole moments, as outlined in SM Section V C [60]. We do not compute the position-resolved profile of magnetic dipole contributions, but only the overall contribution to the SFG signal, which can be extracted from a simulation of a bulk system, according to a fluctuation-dissipation relation presented in SM Section V C [60]. Hence, we omit magnetic dipole contributions, whenever we present the response profiles $\tilde{s}^{(2)}_{ijk}(z)$, but consider them when we present the total spectra $\tilde{S}^{(2)}_{ijk}$. In contrast to the electric dipole and electric quadrupole contributions, the magnetic dipole contribution does depend on $\omega^{\mathrm{VIS}}$}, set to $2730\text{\,}\mathrm{T}\mathrm{H}\mathrm{z}$, which is the center frequency of the VIS-field employed in the experimental measurement of the bending band [50] we compare with. The VIS frequency in the experimental reference SFG spectra of the stretching band [31, 117] is quite similar.
Appendix D Linear Dielectric and Absorption Profiles
We define the linear response of the polarization density $p_{i}(z,t)=D_{i}(z,t)-\varepsilon_{0}E_{i}(z,t)$} to an external field of amplitude $\mathcal{F}^{\alpha}_{i}$ as
$\varepsilon_{0}^{-1}p^{(1)}_{i}(z,t)=e^{-i\omega^{\alpha}t}\tilde{s}^{(1,\mathrm{P})}_{ij}(z)\mathcal{F}_{j}^{\alpha}+c.c.\,.$ (23)
The extraction of $\tilde{s}^{(1,\mathrm{P})}_{ij}(z)$ from molecular dynamics simulation has been described before [80, 81, 82, 83, 84, 85, 86] and is reproduced in SM Sections V B [60]. At the interface, the dielectric profile is tensorial, and the component perpendicular to the interface is given by
$\tilde{\varepsilon}_{zz}^{\mathrm{\alpha}}(z)=\frac{\varepsilon_{0}^{-1}\mathcal{D}^{\alpha}_{z}}{\mathcal{E}^{\alpha}_{z}(z)}=\frac{1}{1-\tilde{s}^{(1,\mathrm{P})}_{zz}(z)}\,$, (24)
while the component parallel to the interface is determined by
$\tilde{\varepsilon}^{\alpha}_{xx}(z)=\frac{\varepsilon_{0}^{-1}\mathcal{D}^{\alpha}_{x}(z)}{\mathcal{E}^{\alpha}_{x}}=1+\tilde{s}^{(1,\mathrm{P})}_{xx}(z).$ (25)
Here, $\mathcal{D}^{\alpha}_{i}(z)$ and $\mathcal{E}^{\alpha}_{i}(z)$, are defined in Equations (10) and (9), respectively. Due to the symmetry of our system we have $\tilde{\varepsilon}^{\alpha}_{yy}(z)=\tilde{\varepsilon}^{\alpha}_{xx}(z)$. In bulk, the dielectric tensor reduces to the isotropic component
$\tilde{\varepsilon}^{\alpha}(z)=\frac{\tilde{\varepsilon}^{\alpha}_{xx}(z)+\tilde{\varepsilon}^{\alpha}_{yy}(z)+\tilde{\varepsilon}^{\alpha}_{zz}(z)}{3}\,.$ (26)
Further information about the dielectric profiles is given in SM Section XI [60].
Appendix E Electric Dipole SFG Contribution as a Fingerprint for Interfacial Structure
The response of the second-order electric current density to spatially constant external fields, which are $z$-polarized D-fields and $x$ or $y$-polarized E-fields, is given in Equation (7). As we want to relate SFG spectra to the molecular orientation distribution, we are interested in the second-order response of the electric dipole density to the average amplitude of the local E-field $\mathcal{E}^{\mathrm{L},\alpha}_{i}(z)$ acting on the molecular centers. We define the second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ by
$\varepsilon_{0}^{-1}j_{i}^{(2,\mathrm{DD})}(z,t)=\\ -i\omega^{\mathrm{SFG}}f^{\mathrm{SFG}}_{i}(z)e^{-i\omega^{\mathrm{SFG}}t}\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)\mathcal{E}^{\mathrm{L,VIS}}_{j}(z)\mathcal{E}^{\mathrm{L,IR}}_{k}(z)+c.c.\,.$ (27)
The second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ plays a central role in SFG theory as it allows to relate the macroscopic electric dipole contribution to the molecular orientation [27, 25, 28, 29]. We can relate $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ to the second-order response function by
$\tilde{s}^{(2,\mathrm{DD})}_{ijk}\left(z\right)=f^{\mathrm{SFG}}_{i}(z)f^{\mathrm{VIS}}_{j}(z)f^{\mathrm{IR}}_{k}(z)\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)\,$, (28)
where $f^{\alpha}_{i}(z)$ are the laterally averaged local field factors, defined in Equation (1). This is the formalism traditionally used in SFG theory [68, 41, 24, 25, 31, 42]. Note that we do not introduce any locality approximations in the computation of the second-order response $\tilde{s}^{(2)}_{ijk}(z)$, but rather predict $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ by dividing the nonlocal second-order response profile $\tilde{s}^{(2,\mathrm{DD})}_{ijk}\left(z\right)$ by the averaged local field factors $f^{\alpha}_{i}(z)$. We neglect the frequency dependence of $f^{\alpha}_{i}(z)$ by approximating $f^{\alpha}_{i}(z)\approx f^{\mathrm{VIS}}_{i}(z)$. We compare $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ with $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}\left(z\right)$, which is solely based on the anisotropic orientation of the molecules.
Figure 6: Sketch of the molecular Eckart frame for a water molecule.
The Euler angles $\phi$, $\theta$ and $\psi$ specify the molecule’s orientation and we employ the $z^{\prime}y^{\prime}z^{\prime}$ convention [118]. We define the molecular Eckart frame [110, 111, 119, 120] by aligning the molecular $z^{\prime}$-axis with the permanent dipole vector. The molecular $x^{\prime}$-axis points out of the molecular plane, while the molecular $y^{\prime}$-axis is chosen to be perpendicular to both $x^{\prime}$ and $z^{\prime}$. The molecular Eckart frame is depicted in Figure 6. As shown in SM Section VII [60], the molecular hyperpolarizability $\tilde{\beta}_{ijk}(\theta,\psi)$ is determined by three orientation-dependent functions $q_{10}(\theta)$, $q_{30}(\theta)$ and $q_{32}(\theta,\psi)$. The first and third Legendre polynomials [118]
$q_{10}(\theta)$ $=\cos{\theta}$ (29)
$q_{30}(\theta)$ $=\frac{1}{2}\left[5\cos^{3}{\theta}-3\cos{\theta}\right]\,$ (30)
describe the orientation of the molecular dipoles, while $q_{32}(\theta,\psi)$ accounts for the rotation of the molecule around the molecular dipole moment and is only relevant if the molecule is not uniaxial. Consequently, $q_{32}(\theta,\psi)$ is the biaxiality parameter, given by
$q_{32}(\theta,\psi)$ $=\cos(\theta)(\sin^{2}{\theta}\cos^{2}{\psi}-\sin^{2}{\theta}\sin^{2}{\psi})\,.$ (31)
We obtain $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ by spatial averaging all hyperpolarizabilities according to
$\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ $=\frac{1}{V}\sum\limits_{n}^{N_{\mathrm{mol}}}\tilde{\beta}_{ijk}(\theta^{n},\psi^{n})$ (32)
$=\rho(z)\left[q_{10}(z)\tilde{\beta}_{ijk}^{10}+q_{30}(z)\tilde{\beta}_{ijk}^{30}+q_{32}(z)\tilde{\beta}_{ijk}^{32}\right]\,$, (33)
where $V$ is the volume, $\rho(z)$ is the molecular density and $q_{lm}(z)$ is the z-dependent average of $q_{lm}(\theta^{n},\psi^{n})$ over all molecules. The frequency dependent coefficients $\tilde{\beta}_{ijk}^{lm}$ are presented in Figure 3 (j). A detailed description of the orientation analysis can be found in SM Section VII [60]. The full information about the orientation distribution is determined by the orientation distribution function defined by
$\rho_{\mathrm{ORI}}(\theta,\psi)=\frac{1}{2\pi\sin{\theta}N_{\mathrm{mol}}}\sum\limits_{n}^{N_{\mathrm{mol}}}\left\langle\delta(\theta-\theta^{n})\delta(\psi-\psi^{n})\right\rangle\,.$ (34)
The prefactor ensures that $\rho_{\mathrm{ORI}}(\theta,\psi)$ is properly normalized as $\int\limits_{0}^{\mathrm{2\pi}}\mathrm{d}\phi\int\limits_{0}^{\mathrm{\pi}}\sin{\theta}\mathrm{d}\theta\int\limits_{0}^{\mathrm{2\pi}}\mathrm{d}\psi\,\rho_{\mathrm{ORI}}(\theta,\psi)=1$. In an isotropic system $\rho_{\mathrm{ORI}}(\theta,\psi)$ is constant and is given by $\frac{1}{8\pi^{2}}$.